# Deciphering black hole phase transitions through photon spheres

### v1

@@ Abstract
Black hole thermodynamics is a crucial and foundational aspect of black hole physics, yet its observational verification remains exceptionally challenging. The photon sphere of a black hole, a manifestation of strong gravitational effects, is intrinsically linked to its shadow, which has been directly captured through observations made by the Event Horizon Telescope. Investigating black hole thermodynamics from a gravitational perspective presents an intriguing avenue for research. This paper investigates the relationship between the thermodynamic phase transition and the photon sphere of a black hole with quantum anomaly. It proposes that the photon sphere encodes information about the black hole phase transition, arguing that the change in the photon sphere radius can serve as an order parameter characterizing the black hole’s phase transition.
@@ I Introduction
The photon sphere of a black hole, a manifestation of strong gravitational effects, is intrinsically linked to its shadow, as directly observed by the Event Horizon Telescope. Investigating black hole thermodynamics from a gravitational perspective is an intriguing subject. Inspired by this perspective, we extend our team’s prior work by further exploring the relationship between black hole thermodynamic phase transitions and the photon sphere in the context of black holes exhibiting quantum anomalies. These anomalies violate conventional thermodynamic scaling laws, with first-order phase transitions occurring both below and above the critical point.
@@ II.2 Critical points and coexistence curve
In the extended phase space, the thermodynamic behavior of the charged black hole with trace anomaly exhibits a rich and intricate phase structure. As will be shown later, the system displays swallowtail behavior and features two critical points. In this subsection, we focus on analyzing the phase structure and deriving the coexistence curve.
@@
From the above equations, we can get two critical points for $80\leq\alpha_{\text{c}}\leq Q^{2}/$ 8. These critical points are given by
@@
where we have defined $\Gamma=\sqrt{192\alpha_{\text{c}}^{2}-96\alpha_{\text{c}}Q^{2}+9Q^{4}}$. For $\alpha_{\text{c}}=Q^{2}/8$, the thermodynamic behavior of the black hole with quantum anomaly exhibits unexpected phase transitions, with critical exponents deviating from conventional ones and violating scaling laws [66]. Under the condition $\alpha_{\text{c}}=Q^{2}/8$, the two previously distinct critical points (15) and (16) merge into a single one, which is
@@
In the following, we consider the case $\alpha_{\text{c}}=Q^{2}/8$.
@@
To investigate the coexistence curve in the $P-V$ plane, we consider the Gibbs free energy in the canonical ensemble with fixed temperature $T$ and charge $Q$. We depict the Gibbs free energy $G$ as a function of the thermodynamic pressure $P$ in Fig. 1. As shown in the figure, swallowtail behavior occurs both below and above the critical point but disappears precisely at the critical point.
In addition to the swallowtail behavior observed in the Gibbs free energy during the first-order phase transition, oscillatory behavior also appears in a pair of conjugate thermodynamic variables, as illustrated in the $P-V$ plane in Fig. 2. This oscillation disappears at the critical temperature.
Figure 2: The oscillatory curve in the $P-V$ plane. Below and above the critical point, a first-order phase transition takes place, characterized by oscillatory behavior in a pair of conjugate thermodynamic variables. At the critical point, the oscillatory behavior disappears.
in the canonical ensemble with fixed temperature $T$ and charge $Q$, we have
@@
This result shows that the areas of the shadowed regions in Fig. 3 are equal. This suggests that Maxwell’s equal law holds in the $P-V$ plane.
@@
From Maxwell’s equal area law in the $P-V$ plane, Eq. (22), we have
@@
Figure 3: Maxwell’s equal area law in the $P-V$ plane. The areas in the two shadowed regions are the same. Here we have chosen $T=1.3T_{\text{c}}$.
@@ III.1 Photon sphere for black hole with trace anomaly
where $E$ and $L$ are the energy and angular momentum of the photon, respectively. The radial momentum of the photon is
$P_{r}=\frac{\dot{r}}{f(r)}.$ (33)
From the energy and angular momentum of the photon, we can get the $t$ motion and $\phi$ motion, which are
$\dot{t}$ $=\frac{E}{f(r)},$ (34)
$\dot{\phi}$ $=\frac{L}{r^{2}}.$ (35)
@@
$02f(r_{\text{ps}})-r_{\text{ps}}\partial_{r}f(r_{\text{ps}})=$ 0 . (40)
@@
At the critical point of the black hole phase transition, the photon sphere radius is
$r_{\text{phc}}=2.35788Q.$ (42)
To explore the relationship between the photon sphere and the thermodynamic phase transition, we define the reduced photon sphere radius
$\tilde{r}_{\text{ps}}=\frac{r_{\text{ps}}}{r_{\text{psc}}}.$ (43)
@@ III.2 Unstable photon sphere and phase transition
As illustrated in Figs. 11 and 2, the Gibbs free energy exhibits a characteristic swallowtail structure, while a pair of the conjugate thermodynamic variables display oscillatory behavior during the small–large black hole phase transition. At the critical point, both the swallowtail structure and the oscillatory behavior disappear. In this subsection, we investigate the characteristic behaviors of the photon sphere during black hole phase transitions and examine the potential correlation between these transitions and the properties of the photon sphere.
@@
Equations (45) and (46) give a parametric function for the reduced photon sphere radius $\tilde{r}_{\text{ps}}$
@@
From Figs. 7 and 7, we can see that there are oscillatory behaviors and there exist two extremal points in the $p-\tilde{r}_{\text{phs}}$ curve for the first-order phase transition, and the oscillatory behavior disappears at the critical point.
@@ III.3 Order parameter and critical exponent
Order parameter is an important quantity to characterize phase transition phenomena. It changes discontinuously in a first-order phase transition but continuously at the critical point. The order parameter goes to zero at the critical point along the coexistence curve. In this subsection, we explore whether we can use quantities related to the photon sphere as an order parameter to depict black hole phase transition.
@@
The reduced photon sphere radius, $r_{\text{ps}}/r_{\text{psc}}$, is plotted as a function of the phase transition temperature, $T/T_{\text{c}}$, in Fig. 8. In the reduced parameter space, the reduced photon radius $\tilde{r}_{\text{ps}}$ and phase transition temperature $\tau$ are found to be independent of charge. As illustrated in the figure, below the critical point, the photon sphere radius $\tilde{r}_{\text{ps}}$ increases for the saturated coexistence small black hole, whereas it decreases for the saturated coexistence large black hole as the temperature increases. The reduced photon radii $\tilde{r}_{\text{ps}}$ for the coexistence small and large black holes are the same at the critical point $\tau=1$. As the temperature increases, the reduced photon radius $\tilde{r}_{\text{ps}}$ increases for the coexisting large black hole, whereas it initially decreases and then increases for the coexisting small black hole.
We also plot the difference of the reduced photon sphere radius for the saturated coexistence large and small black holes $\Delta\tilde{r}_{\text{ps}}$ in Fig. 8. From the figure, we can see that the difference of the reduced photon radius decreases with the transition temperature and goes to zero at the critical point. In contrast to the behavior observed in the Reissner-Nordström AdS black hole [56], where the difference of the photon sphere radius for the coexistence of large and small black holes terminates at the critical point, the difference of the photon sphere radii for the coexistence of large black holes and small black holes increases with the transition temperature when the temperature exceeds the critical temperature.
The behavior of the difference of the reduced photon sphere radius $\Delta\tilde{r}_{\text{ps}}$ is extremely similar to the behavior of the order parameter. It exhibits nonzero values during the first-order phase transition and vanishes at the critical point. But we still need to check the critical exponent of the parameter. Near the critical point, we have
@@
From the definition of the order parameter, it is clear that the changes of the reduced photon sphere radius for saturated coexistence large and small black holes can be regarded as the order parameter. As suggested in Ref. [66], the presence of a quantum anomaly leads to a critical exponent of $\beta=1$ for the order parameter, which markedly differs from the value $\beta=1/2$ observed in the Reissner-Nordström AdS black hole and in most other known black hole cases. Furthermore, the quantum anomaly causes the critical exponents to violate the scaling laws predicted by mean field theory [66].
@@ IV Conclusion and discussion
In this study, we examined the relationship between the thermodynamic phase transition and the photon sphere of a static spherically symmetric charged black hole with trace anomaly. The charged black hole with trace anomaly can be regarded as a quantum correction of the Reissner-Nordström AdS black hole by incorporating the backreaction induced by the quantum conformal anomaly of the quantum field. Unlike the Reissner-Nordström AdS black hole, its thermodynamic behavior exhibits two distinct critical points. We considered the special case $\alpha_{\text{c}}=Q^{2}/8$, where the two critical points merge, and the resulting critical exponents deviate from the scaling laws predicted by mean field theory. In the extended phase space, below and above the critical point, the Gibbs free energy exhibits the characteristic swallowtail behavior indicating a first-order phase transition. At the critical point, the swallowtail behavior disappears. Below and above the critical temperature, there exists an oscillatory behavior in the $P-V$ plane. The oscillatory behavior disappears at the critical temperature. From the oscillatory behavior in the $P-V$ plane, we found that Maxwell’s equal area law holds, and we obtained the analytical coexistence curve from Maxwell’s equal area law. Similar to the case of the Reissner-Nordström AdS black hole, the coexistence curve is independent of the electric charge in the reduced parameter space.
@@
After having explored the main thermodynamic features for the charged black hole with trace anomaly, we explored the relationship between the thermodynamical phase transition and the photon sphere. We found that the thermodynamic variables such as the black hole temperature and thermodynamic pressure oscillate with the unstable photon sphere radius and there are two extreme points when there is a first-order phase transition, and the oscillatory behavior disappears at the critical point. This phenomenon strongly hints that there is a relationship between the unstable photon sphere and black hole phase transition. We further explored the changes of the reduced photon sphere radius for the coexistence large and small black holes; we found it goes to zero at the critical point and changes discontinuously in a first-order phase transition. Furthermore, the critical exponent for the changes of the photon sphere radius for the coexistence large and small black holes is $\beta=1$, which is the same as the order parameter obtained in Ref. [66]. This suggests that the changes of the reduced photon sphere radius can be regarded as the order parameter for the small-large black hole phase transition.
@@
For the Reissner-Nordström AdS black hole, the coexistence curve for small-large black holes terminates at the critical point. An early work showed that the change of the photon sphere radius for $d$-dimensional Reissner-Nordström AdS black holes can serve as an order parameter for black hole phase transition [56]. When the backreaction of the quantum conformal anomaly is taken into account, the metric of the Reissner-Nordström AdS black hole is modified and the coexistence curve for small-large black holes does not terminate at the critical point. We further confirmed that the changes of the photon sphere radius for black holes with trace anomaly can also be regarded as the order parameter even though the coexistence curve does not terminate at the critical point. Together with investigations for the relationship between the photon sphere and phase transitions for other black holes [57, 56, 58], our work further enhances that the photon sphere might encode information about black hole phase transitions.
@@
These investigations provide a promising avenue for exploring thermodynamic effects from a gravitational perspective, indicating that black hole shadows may encode valuable thermodynamic information. In addition to shadows, other observable phenomena—such as particle motion and chaotic dynamics—may also serve as effective probes of black hole phase transitions and might be worth further investigation [70, 71].

### v2

@@ Abstract
Black hole thermodynamics is a crucial and foundational aspect of black hole physics, yet its observational verification remains exceptionally challenging. The photon sphere of a black hole, a manifestation of strong gravitational effects, is intrinsically linked to its shadow, which has been directly captured through observations made by the Event Horizon Telescope. Investigating black hole thermodynamics from a gravitational perspective presents an intriguing avenue for research. This paper obtains an analytical formula for the coexistence curve and investigates the relationship between the thermodynamic phase transition and the photon sphere of a black hole with quantum anomaly. It proposes that the photon sphere encodes information about the black hole phase transition, arguing that the change in the photon sphere radius can serve as an order parameter characterizing the black hole’s phase transition.
@@ I Introduction
The photon sphere of a black hole, a manifestation of strong gravitational effects, is intrinsically linked to its shadow, as directly observed by the Event Horizon Telescope. Investigating black hole thermodynamics from a gravitational perspective remains a compelling subject to research. Inspired by this perspective, we further explore the relationship between black hole thermodynamic phase transitions and the photon sphere in the presence of quantum anomaly. This anomaly intrinsically modifies the scaling behavior and leads to a breakdown of conventional scaling relations, with first-order phase transitions occurring both below and above the critical point.
@@ II.2 Critical points and coexistence curve
In the extended phase space, the thermodynamic behavior of the charged black hole with trace anomaly exhibits a rich and intricate phase structure. As will be shown later, the system displays swallowtail behavior and features two critical points. In this subsection, we focus on analyzing the phase structure and deriving the coexistence curve. The main new result presented in this subsection is the derivation of the analytical expression for the coexistence curve, which, remarkably, is found to be independent of the black hole charge in the reduced parameter space. While in previous work the coexistence curve was obtained numerically via point-by-point plotting [66], our work provides a fully analytical form.
@@
From the above equations, we can get two critical points for $0\leq\alpha_{\text{c}}\leq Q^{2}/8$. These critical points are given by
@@
where we have defined $\Gamma=\sqrt{192\alpha_{\text{c}}^{2}-96\alpha_{\text{c}}Q^{2}+9Q^{4}}$. For $0\leq\alpha_{\text{c}}<Q^{2}/8$, there are two critical points, and the critical exponents satisfy the scaling law. However, when $\alpha_{\text{c}}=Q^{2}/8$, the two distinct critical points (15) and (16) merge into a single one, which is
@@
In this case, the thermodynamic behavior of black holes with quantum anomalies reveals novel and unexpected phase transitions. The Gibbs free energy exhibits a characteristic swallowtail structure not only below but also above the critical pressure, and critical exponents deviate from the conventional values and violate standard scaling laws. These deviations arise intrinsically from the quantum conformal anomaly, which fundamentally alters the scaling relations and establishes a clear causal link between quantum effects and scaling law violation [66]. In the following, we consider the case $\alpha_{\text{c}}=Q^{2}/8$.
@@
To investigate the coexistence curve in the PVP-V plane, we consider the Gibbs free energy in the canonical ensemble with fixed temperature $T$ and charge $Q$. We depict the Gibbs free energy $G$ as a function of the thermodynamic pressure $P$ in Fig. 1. As shown in the figure, swallowtail behavior occurs both below and above the critical point but disappears precisely at the critical point.
At the phase transition point in the canonical ensemble with fixed temperature $T$ and charge $Q$, the Gibbs free energies of the small and large black hole phases become equal. Consequently, the change in Gibbs free energy across the transition satisfies
@@
This result shows that the areas of the shadowed regions in Fig. 2 are equal. This suggests that Maxwell’s equal law holds in the PVP-V plane.
@@
From Maxwell’s equal area law in the PVP-V plane, Eq. (22), we have
@@
Figure 2: Maxwell’s equal area law in the PVP-V plane. The curve oscillates for a pair of conjugate thermodynamic variables as a first-order phase transition takes place and the areas in the two shadowed regions are the same. Here we have chosen $T=1.3T_{\text{c}}$.
@@ III.1 Photon sphere for black hole with trace anomaly
where $E$ and $L$ are the energy and angular momentum of the photon, respectively.
@@
$2f(r_{\text{ps}})-r_{\text{ps}}\partial_{r}f(r_{\text{ps}})=0.$ (37)
@@
The impact parameter, which determines the apparent size of the black hole shadow, is related to its radius through the formula [70]
$r_{\text{sh}}=r_{\text{ps}}\sqrt{\frac{f(r_{0})}{f(r_{\text{ps}})}}=b_{\text{ps}}\sqrt{f(r_{0})},$ (39)
here, $r_{0}$ denotes the position of the observer. The result indicates that the radius of the black hole shadow is proportional to the impact parameter associated with the photon sphere. Therefore, for convenience, we use the impact parameter as a measure of the shadow radius.
At the critical point of the black hole phase transition, the radius of the photon sphere and the corresponding impact parameter are found to be
$r_{\text{phc}}$ $=2.35788Q,$ (40)
$b_{\text{phc}}$ $=2.08213Q.$ (41)
Remarkably, the critical impact parameter is smaller than the critical photon sphere radius, which stands in sharp contrast to the case of massive black holes in asymptotically flat spacetimes.
To explore the relationship between the photon sphere and the thermodynamic phase transition, we define the reduced photon sphere radius and reduced impact parameter
$\tilde{r}_{\text{ps}}$ $=\frac{r_{\text{ps}}}{r_{\text{psc}}},$ (42)
$\tilde{b}_{\text{ps}}$ $=\frac{b_{\text{ps}}}{b_{\text{psc}}}.$ (43)
@@ III.2 Unstable photon sphere and phase transition
As illustrated in Figs. 11 and 2, the Gibbs free energy exhibits a characteristic swallowtail structure, while a pair of the conjugate thermodynamic variables displays oscillatory behavior during the small–large black hole phase transition. At the critical point, both the swallowtail structure and the oscillatory behavior disappear. In this subsection, we investigate the characteristic behaviors of the photon sphere during black hole phase transitions and examine the potential correlation between these transitions and the properties of the photon sphere.
@@
Equations (45) and (46) give a parametric function for the reduced photon sphere radius $\tilde{r}_{\text{ps}}$ .
@@
From Figs. 6 and 6, we can see that there are oscillatory behaviors and there exist two extremal points in the $p-\tilde{r}_{\text{phs}}$ curve for the first-order phase transition, and the oscillatory behavior disappears at the critical point.
@@ III.3 Order parameter and critical exponent
Order parameter is an important quantity to characterize phase transition phenomena. It changes discontinuously in a first-order phase transition but continuously at the critical point. The order parameter goes to zero at the critical point along the coexistence curve. In this subsection, we explore whether we can use quantities related to the photon sphere as an order parameter to depict black hole phase transitions.
@@
Figure 8: (a). The reduced impact parameter for the coexistence small and large black holes. (b). The difference of the reduced impact parameter for the coexistence large and small black holes.
The reduced photon sphere radius, $r_{\text{ps}}/r_{\text{psc}}$ and the reduced impact parameter $b_{\text{ps}}/b_{\text{psc}}$ , are plotted as a function of the phase transition temperature, $T/T_{\text{c}}$, in Figs. 7 and 8. In the reduced parameter space, the reduced photon radius $\tilde{r}_{\text{ps}}$, the reduced impact parameter $\tilde{b}_{\text{ps}}$, and phase transition temperature $\tau$ are found to be independent of charge. As illustrated in the figure, below the critical point, the photon sphere radius $\tilde{r}_{\text{ps}}$ increases for the saturated coexistence small black hole, whereas it decreases for the saturated coexistence large black hole as the temperature increases. The reduced photon radii $\tilde{r}_{\text{ps}}$ for the coexistence small and large black holes are the same at the critical point $\tau=1$. As the temperature increases, the reduced photon radius $\tilde{r}_{\text{ps}}$ increases for the coexisting large black hole, whereas it initially decreases and then increases for the coexisting small black hole.
We also plot the difference of the reduced photon sphere radius and the difference of the reduced impact parameter for the saturated coexistence large and small black holes in Figs. 7 and 8. From the figure, we can see that the difference of the reduced photon radius decreases with the transition temperature and goes to zero at the critical point. In contrast to the behavior observed in the Reissner-Nordström AdS black hole [56], where the difference of the photon sphere radius for the coexistence of large and small black holes terminates at the critical point, the difference of the photon sphere radii for the coexistence of large black holes and small black holes increases with the transition temperature when the temperature exceeds the critical temperature.
The behavior of the difference of the reduced photon sphere radius $\Delta\tilde{r}_{\text{ps}}$ is extremely similar to the behavior of the order parameter. It exhibits nonzero values during the first-order phase transition and vanishes at the critical point. But we still need to check the critical exponent of the parameter. By performing a series expansion of the reduced photon sphere radius near the critical point, we obtain an analytical expression
@@
The impact parameter is directly related to the radius of black hole shadows. Similarly, near the critical point, the difference of the reduced impact parameter for large and small black hole phases can be expressed as
$\begin{split}\Delta\tilde{b}_{\text{ps}}\simeq\left(\frac{\partial\tilde{b}_{\text{ps}}}{\partial r_{\text{h}}}\right)_{\text{c}}\left(r_{\text{hl}}-r_{\text{hs}}\right)\simeq 0.46\mid\tau-1\mid.\end{split}$ (49)
The critical exponent for the reduced impact parameter is still $\beta=1$.
From the definition of the order parameter, it is clear that the changes of the reduced photon sphere radius and the reduced impact parameter for saturated coexistence large and small black holes can be regarded as the order parameter. As suggested in Ref. [66], the presence of a quantum anomaly leads to a critical exponent of $\beta=1$ for the order parameter, which markedly differs from the value $\beta=1/2$ observed in the Reissner-Nordström AdS black hole and in most other known black hole cases. Furthermore, the quantum anomaly causes the critical exponents to violate the scaling laws predicted by mean field theory [66].
@@ IV Conclusion and discussion
In this study, we examined the relationship between the thermodynamic phase transition and the photon sphere of a static spherically symmetric charged black hole with trace anomaly. The charged black hole with trace anomaly can be regarded as a quantum correction of the Reissner-Nordström AdS black hole by incorporating the backreaction induced by the quantum conformal anomaly of the quantum field. Unlike the Reissner-Nordström AdS black hole, its thermodynamic behavior exhibits two distinct critical points. We considered the special case $\alpha_{\text{c}}=Q^{2}/8$, where the two critical points merge, and the resulting critical exponents deviate from the scaling laws predicted by mean field theory. In the extended phase space, below and above the critical point, the Gibbs free energy exhibits the characteristic swallowtail behavior indicating a first-order phase transition. At the critical point, the swallowtail behavior disappears. Below and above the critical temperature, there exists an oscillatory behavior in the PVP-V plane. The oscillatory behavior disappears at the critical temperature. From the oscillatory behavior in the PVP-V plane, we found that Maxwell’s equal area law holds, and we obtained the analytical coexistence curve from Maxwell’s equal area law. Similar to the case of the Reissner-Nordström AdS black hole, the coexistence curve is independent of the electric charge in the reduced parameter space.
@@
After having explored the main thermodynamic features for the charged black hole with trace anomaly, we explored the relationship between the thermodynamical phase transition and the photon sphere. We found that the thermodynamic variables such as the black hole temperature and thermodynamic pressure oscillate with the unstable photon sphere radius and there are two extreme points when there is a first-order phase transition, and the oscillatory behavior disappears at the critical point. This phenomenon strongly hints that there is a relationship between the unstable photon sphere and black hole phase transition. We further explored the changes of the reduced photon sphere radius for the coexistence large and small black holes; we found it goes to zero at the critical point and changes discontinuously in a first-order phase transition. Furthermore, the critical exponent for the changes of the photon sphere radius for the coexistence large and small black holes is $\beta=1$, which is the same as the order parameter obtained in Ref. [66]. This suggests that the changes of the reduced photon sphere radius can be regarded as the order parameter for the small-large black hole phase transition. In the reduced parameter phase space, all the results for the relationship between phase transitions and the photon sphere are obtained analytically and they are independent of black hole charge.
@@
For the Reissner-Nordström AdS black hole, the coexistence curve for small-large black holes terminates at the critical point. An early work showed that the change of the photon sphere radius for $d$-dimensional Reissner-Nordström AdS black holes can serve as an order parameter for black hole phase transition [56]. When the backreaction of the quantum conformal anomaly is taken into account, the metric of the Reissner-Nordström AdS black hole is modified and the coexistence curve for small-large black holes does not terminate at the critical point. We further confirmed that the changes of the photon sphere radius for black holes with trace anomaly can also be regarded as the order parameter even though the coexistence curve does not terminate at the critical point.
The photon sphere encodes geometric and optical information [71, 72], which is directly linked to observable features such as the black hole shadow. By choosing the photon sphere radius as the order parameter, our aim is to bridge thermodynamic phase transitions and astrophysical observables in a more direct and meaningful way. In this work, we extend the study to phase transitions involving isolated critical points, and further demonstrate that the photon sphere can effectively encode critical behavior. Our results highlight how such thermodynamic phenomena may be imprinted in the photon sphere structure, which could potentially be tested observationally.
As astrophysical black holes are neutron and rotating, for a rotating AdS black hole, thermodynamic information might be effectively encoded in the photon ring, which corresponds to unstable null geodesics confined to the equatorial plane. The properties of the photon ring are determined by the horizon radius, angular momentum, and pressure. Interestingly, it exhibits behavior analogous to that of the photon sphere in spherically symmetric AdS black holes [57], suggesting a broader universality. In this work, we explored whether such correspondence persists in the presence of quantum anomaly, which intrinsically modifies the scaling relations and leads to the violation of conventional scaling law. Together with investigations for the relationship between the photon sphere and phase transitions for other black holes [57, 56, 58], our work further enhances that the photon sphere might encode information about black hole phase transitions.
Although our universe is asymptotically de Sitter, much of the research on thermodynamics for black holes is carried out in asymptotically AdS. The primary reason lies in the well-defined thermodynamic variables and rich phase behaviors for black holes in AdS space. In AdS space, the gravitational potential of AdS space acts like a confining box, allowing black holes to reach thermal equilibrium [15]. Moreover, asymptotically AdS black holes provide a gauge dual description of finite-temperature CFTs through the AdS/CFT correspondence [12]. However, black hole thermodynamics in dS space is more intricate due to two factors: the presence of both cosmological and event horizons places the system between them in a nonequilibrium state, and the lack of a globally timelike Killing vector hinders the definition of asymptotic mass [36]. Our exploration of AdS black hole thermodynamics through the photon sphere may shed light on future studies of de Sitter black hole thermodynamics.
@@
These investigations provide a promising avenue for exploring thermodynamic effects from a gravitational perspective, indicating that black hole shadows may encode valuable thermodynamic information. In addition to shadows [73, 74], other observable phenomena—such as particle motion and chaotic dynamics—may also serve as effective probes of black hole phase transitions and might be worth further investigation [75, 76, 77].